# Forecasting the U.S. Renewable-Energy Mix with an ALR-BDARMA Compositional Time-Series Framework

### v3

@@ 6 Conclusion
This study develops and tests a Bayesian Dirichlet ARMA model whose additive‑log‑ratio mean follows a parsimonious VAR(2) while the Dirichlet concentration parameter evolves with an intercept plus ten monthly harmonics. In a 61‑split rolling evaluation the Dirichlet‑based approach delivers the sharpest and best‑calibrated twelve‑month density forecasts for the U.S. renewable‑energy mix. Relative to a Gaussian VAR(2) with identical seasonal dummies, mean CRPS falls by roughly one‑quarter at short leads and by one‑fifth at a full year; gains over both a seasonal naïve and an additive‑log‑ratio random walk are still larger. Point accuracy measured by Aitchison RMSE is on par with the VAR through eight months and remains within 0.02 units thereafter, underscoring that the improvement is driven by better uncertainty quantification rather than a sacrifice in central tendency.
The single‑origin forecast issued for calendar year 2025 illustrates practical implications. BDARMA produces calibrated 90% intervals that tighten for low‑volatility sources such as geothermal and waste while remaining wide enough for wind and solar to capture intra‑year swings. Because the competing models offer medians only, planners relying on them must add ad‑hoc bands or scenario grids that may violate simplex coherence. In contrast, the Dirichlet framework provides a one‑stop forecast that can feed directly into capacity‑expansion studies, transmission planning, and reliability assessments, all of which depend critically on both the mean mix and its dispersion.
Methodologically the model is straightforward to extend. Exogenous drivers like fuel prices, policy shocks, regional weather indices—can be added as additional columns in the mean or precision design matrices, and the hierarchical structure naturally scales to state or utility sub‑regions. Future work may also explore coupling the Dirichlet layer with non‑linear latent dynamics or integrating the forecast densities into stochastic power‑system optimization. For the present, the results demonstrate that a modest amount of Bayesian structure, applied in the right geometric space, is enough to yield substantial predictive gains and actionable insights for the energy‑transition debate.

### v4

@@ 6 Conclusion
This paper develops and evaluates a Bayesian Dirichlet ARMA (BDARMA) model for monthly U.S. renewable‑energy shares, with a VAR(2) mean in additive‑log‑ratio space and a seasonal Dirichlet precision. In a 61‑split rolling evaluation, BDARMA delivers the sharpest, best‑calibrated twelve‑month density forecasts: mean CRPS is lower than a transform‑space VAR(2) at every horizon and far below naïve rules, while point accuracy (Aitchison RMSE) matches the VAR through eight months and stays within roughly two hundredths thereafter. The result is improved uncertainty quantification without sacrificing the central path.
The errors have planning implications by technology. Hydro and wind, our most seasonal components, retain wider, seasonally patterned bands that inform spring runoff scheduling and winter ramping reserves. Solar’s trend plus summer plateau yields medium‑horizon gains that tighten estimates of midday surplus risk and storage needs. Geothermal and waste are near‑deterministic, justifying narrow tolerance bands, while wood’s broader dispersion argues for conservative procurement where fuel and policy are more volatile.
A balanced view also clarifies when simpler models are reasonable. VAR(2) can be preferred when decisions hinge almost entirely on point forecasts, exogenous regressors are abundant, and computational speed is paramount; the seasonal naïve is defensible for steady systems and as a monitoring baseline. By contrast, BDARMA should be the default when calibrated densities, coherence across components, and explicit seasonality matter for reserves, transmission, and storage planning.